OpenStudy (anonymous):

how to solve this using absolute value... /X^2-3X-10/=5

OpenStudy (anonymous):

|x^2 - 3x - 10| = 5 x^2 - 3x -10 = 5 and x^2 -3x-10 = -5

OpenStudy (anonymous):

x^2 - 3x -15 = 0 and x^2 -3x -5 = 0 now can you find x in both the cases

OpenStudy (anonymous):